# Supersonic shear waves in dipolar supersolids

### v3

@@ Abstract
Dipolar supersolids—quantum states that are simultaneously superfluid and solid—have had their superfluid nature rigorously tested, while its solid nature remains uncharted. Arguably, the defining characteristic of a solid is the existence of elastic shear waves. In this work, we investigate transverse wave packet propagation in dipolar supersolids with triangular and honeycomb structure. Remarkably, the honeycomb supersolid displays anomalous dispersion, supporting waves traveling faster than the transverse speed of sound: a supersonic shear wave. While the distinction between anomalous and normal dispersion appears subtle, it has a profound influence on both the transport and scattering properties. For both supersolid phases, we calculate the shear modulus, a key parameter that quantifies the material’s rigidity. Our findings are pertinent to current experimental efforts scrutinizing the fundamental properties of supersolids.
@@
Here, we propose a protocol to excite shear waves in two-dimensional dipolar supersolids, and measure their resulting propagation velocities. We investigate this protocol for triangular and honeycomb lattice structures, which are ground state solutions for a system of ultracold dipolar atoms under planar confinement [41, 42, 43]. Our results reveal distinct dispersive phenomena between the triangular and honeycomb phases, which we probe through the current density dynamic structure factor (DSF) and the Bogoliubov de-Gennes (BdG) method. These investigations show that the honeycomb structure can support a refractive index smaller than unity, thus supporting shear waves traveling faster than the transverse speed of sound. The idea of a supersonic sound mode in classical wave theory can be traced back to Sommerfeld and Brillouin [44, 45, 46, 47], while in the quantum realm, it is analogous to the anomalous phonon dispersion relation observed in superfluid helium-4 [48, 49, 50, 23]. Finally, we find the relationship between the wave packet propagation velocities and the transverse speed of sound, and estimate the shear modulus of dipolar supersolids.
@@
where the dimensionally reduced wave function $\psi(\bm{\rho},t)$ is defined over the infinite plane, $\bm{\rho}=(x,y)$. The 2D interaction and quantum fluctuation terms, $\Phi_{\sigma}$ and $g_{\text{QF}}$, depend on the variational width $\sigma$ in the perpendicular ($z$) direction. We fix the areal number density $n_{\text{2D}}$, and perpendicular harmonic trap with frequency $\omega_{z}=2\pi\times 72.4\mathrm{\,Hz}$. This model and our numerical methods are detailed in the Appendix.
@@
where $u_{\nu}$ and $v_{\nu}$ are the Bogoliubov modes with band index $\nu$ which obey the translational symmetry of the supersolid state. By substituting the perturbed wavefunction into Eq. (2) and linearizing, we obtain the familiar Bogoliubov-de Gennes (BdG) equations [65, 66] (See Appendix C for details). We solve the BdG equations using a Bloch wave expansion of the Bogoliubov modes to obtain the frequencies, $\omega_{\nu}(\bm{k})$, and overlay the three lowest bands for comparison in Fig. 1(c). Note that the near-identical shape of the superfluid and transverse modes is a coincidence of the chosen parameters, and their speeds of sound are distinct in other regions of parameter space [66]. We can extract three characteristic velocities from the BdG results for the transverse branch: 1) the transverse speed of sound $c_{t}=\lim_{k_{y}\rightarrow 0}\omega_{t}/k_{y}$, 2) the phase velocity $v_{p}(k_{y})=\omega_{t}/k_{y}$, which defines the velocity of a wave with given $k_{y}$, and 3) the group velocity, $v_{g}(k_{y})=\partial\omega_{t}/\partial k_{y}$, which defines the velocity of the wave packet envelope and thus the velocity of energy propagation in the medium  [^2] We extract these quantities from the BdG calculations by fitting the transverse branch with a fourth-order polynomial.. For the triangular supersolid, both $v_{g}$ and $v_{p}$ monotonically decrease from $c_{t}$ with increasing wave number, and subsequently the refractive index, $n=c_{t}/v_{p}$, increases from unity with $k_{y}$. This behavior is referred to as normal dispersion and is a hallmark of conventional crystalline solids [68].
@@
Supersonic shear waves in honeycomb supersolids—Applying the same protocol to the honeycomb supersolid [Fig. 2(a)], we observe radically distinct behavior. For the triangular supersolid, the wave packet edge propagates at the speed of sound, where $\omega=0$, as indicated by an enduring rightward COM displacement (colored red in Fig. 1(b)). In contrast, the wave packet edge in the honeycomb supersolid oscillates between left and right displacement [Fig. 2(b)]. Thus, the fastest velocity in the excitation has $\omega>0$, and exceeds the speed of sound.
We can pinpoint the origin of this unexpected finding from the current density DSF, the BdG dispersion and the corresponding characteristic velocities, as shown in Figs. 2(c)-(d)  [^3] Note that as the perturbation dislocates the high density region, we excite many branches, but still dominantly transverse crystal modes.. The transverse branch displays a subtle upward concavity for intermediate wave vector ($0.2\lesssim k_{y}L/\pi\lesssim 0.4$), supporting supersonic $v_{g}$ and $v_{p}$ ($>c_{t}$), and a sub-unity refractive index. This is known as anomalous dispersion and has significant implications for transverse phonon decay. The superlinear dispersion allows for Beliaev damping of transverse phonons, where one phonon spontaneously decays into two [70, 71]. Though rare, similar phenomena has been observed in other systems such as crystalline materials [72, 73], Wigner and Coulomb crystals [74, 75, 76] and superfluid helium-4 [49], the latter being the subject of thirty years of intense experimental and theoretical debate [23]. We speculate that the physical origin of this anomalous dispersion is related to the recently reported shear instability in the honeycomb supersolid [77], and signals the proximity to a transverse branch instability. We elucidate on this interpretation in the End Matter, however this will be substantiated in a subsequent work. To our knowledge, the honeycomb supersolid constitutes the first quantum solid to display anomalous dispersion, the main result of this Letter.
@@
Consider the experimental implementation of this protocol for the triangular supersolid. The edge velocity matches closely with our extracted $c_{t}$ [Fig. 3(a1)], however due to fluctuations such as temperature effects, tracking the edge of the shear wave may be below the fidelity of experiments. Furthermore, in finite size systems the propagation velocity can be greatly reduced at the boundary of the system, as we show in Appendix E. Thus, experimental estimates may drastically underestimate the shear modulus. At much larger densities than shown here, the transverse branch becomes soft, and the shear wave velocity tends to zero. However, this is in the parameter regime where the stripe phase is the ground state solution [42].
The honeycomb state in Fig. 3(a2) shows a supersonic propagation velocity across the parameter space, however this effect is diminished for increasing densities as indicated by the decreasing width of the shaded region. This transition from anomalous to normal dispersion is reminiscent of the behavior seen in superfluid helium-4 as a function of pressure [23]. The anomalous dispersion is maximized at low densities, close to the honeycomb-to-stripe transition, and may be related to the hybridization of modes across this structural phase transition. Refs. [9, 78, 79, 80, 66] define Eq. (5) using the transverse speed of sound, replacing $v_{s}\rightarrow c_{t}$. The prediction of supersonic transverse waves calls this definition into question. Instead, the stiffness is a $k$-dependent quantity, and the shear modulus of the honeycomb supersolid can exceed the one calculated from the speed of sound.
Conclusions—We have demonstrated that dipolar supersolids with two-dimensional crystalline structure are true solids displaying a finite shear modulus. However, the triangular and honeycomb states exhibit shear wave phenomena of a diverging nature. The triangular lattice structure demonstrates normal dispersion, and this modulated superfluid state behaves as an archetypal solid. In contrast, supersolids with honeycomb lattice structure show anomalous dispersion, which supports supersonic shear wave propagation and opens intriguing avenues for future theoretical and experimental exploration. We expect a rich phenomenology to arise in this phase from the novel scattering and spontaneous decay channels of supersonic transverse phonons.
@@
Our work opens up new avenues for non-invasive measurements of dipolar supersolids. In future work, we will explore how defects in the supersolid affect the propagation of shear waves, bringing the ideas of shear wave elastography to the quantum realm. A supersolid can host both solid and superfluid defects, as holes in the crystal structure and quantized vortices, respectively. These superfluid vortices lie in the low-density regions, and are not possible to image directly [38]. How these defects distort and interact with shear waves [95] may provide an indirect measurement of their presence in future dipolar supersolid experiments. Our findings may also provide insights into the propagation of shear waves in the supersolid phase of a neutron star following a starquake or glitch [29].
@@
Appendix C—To validate the dispersion obtained from the current density DSFs, we expand the Bogoliubov modes in Eq. (4) into Bloch waves
$u_{\nu}(\bm{\rho})=u_{\nu,\bm{k}}(\bm{\rho})e^{i\bm{k}\cdot\bm{\rho}},\quad v_{\nu}(\bm{\rho})=v_{\nu,\bm{k}}(\bm{\rho})e^{i\bm{k}\cdot\bm{\rho}}.$ (12)
By substituting the perturbation into the quasi-2D eGPE and linearizing, we obtain the BdG matrix
$\hbar\omega_{\nu}(\bm{k})\begin{bmatrix}u_{\nu,\bm{k}}\\ v_{\nu,\bm{k}}\end{bmatrix}=\begin{bmatrix}\mathcal{L}+\mathcal{M}&-\mathcal{M}\\ \mathcal{M}&-\left(\mathcal{L}+\mathcal{M}\right)\end{bmatrix}\begin{bmatrix}u_{\nu,\bm{k}}\\ v_{\nu,\bm{k}}\end{bmatrix},$ (13)
where the operators are
$\mathcal{L}$ $=-\frac{\hbar^{2}\nabla^{2}}{2m}-\mu+\Phi_{\sigma}+g_{\text{QF}}|\psi|^{3},$ (14)
$\mathcal{M}f$ $=\psi_{0}\mathcal{F}^{-1}\left[\tilde{U}(\bm{k})\mathcal{F}\left[f\psi_{0}^{*}\right]\right]+\frac{3}{2}g_{\mathrm{QF}}|\psi_{0}|^{3}f.$ (15)
We fit the transverse branch obtained from this procedure with a quartic polynomial in order to extract the three characteristic velocities shown in panel (d) in Figs. 1 and 2. In Fig. 6 we show the transverse branches and the quartic fit for the triangular and honeycomb supersolids at $n_{\text{2D}}=2000\,\mu$m-2 and $n_{\text{2D}}=4300\,\mu$m-2 respectively. Similar to Fig. 5, we overlay the linear approximation $\omega=c_{t}k_{y}$, showing that the transverse branch of the honeycomb supersolid exceeds this linear trajectory in Fig. 6(b).
Appendix D—Here we present a simple physical argument for the origin of the anomalous behavior, motivated by the recent work on the shear instability in the honeycomb supersolid of Ref. [77]. In general, the dispersion relation for a gapless branch must have the form $\omega(k)=Ak+Bk^{2}+\mathcal{O}(k^{3})$ for momentum $k$. A phonon instability occurs when the linear part becomes negative, crossing $A=0$. However, in a crystalline system at the Brillouin zone (BZ) edge $\text{d}\omega(k)/\text{d}k=0$. Therefore, near to the shear wave (transverse phonon) instability the band goes as $k^{2}$ for low momenta and flattens at the BZ edge. These criteria set “boundary conditions” for the band and lead to anomalous behavior for intermediate $k$. We note that this implies an enhancement of the anomalous behavior close to the shear wave instability, however there the honeycomb state is not the ground state, and it may not be possible to dynamically probe this in an experiment without entering the stripe regime. Our results show that this behavior persists in an experimentally relevant parameter regime where the honeycomb state is indeed the ground state.
Appendix E—It is pertinent to ask how our results scale in a finite sized system. As the elastic tension for droplets along the edge of the system are only provided by a single neighboring row, can shear waves still propagate in this system? How does the inhomogeneity affect our results? We tackle these questions by considering a system with an additional harmonic trap along one axis, i.e. inserting $V=m\omega_{y}^{2}y^{2}/2$ into our Hamiltonian in Eq. (2), such that our system emulates a segment of a supersolid in a ring trap geometry [99, 100, 101].
[…]
Fig. 7 shows our results. The initial state, in (a), is excited via a vertical shift applied to the outermost column of droplets. We find that, as in the infinite supersolid, a shear wave propagates across the system. Unlike the infinite case, there is now a spatially varying velocity, peaking where the density is largest. We define the local velocity between the columns by the time delay at which pairs of columns reach their maximum displacement. This change in velocity is due to the variation in the inter-droplet spacing between columns. Our results verify that shear waves can propagate in an experimentally accessible geometry. In future work, we will investigate this phenomena in two-dimensional box-traps, where the trap itself will play less of a role [102], or in azimuthally symmetric harmonic traps, where the radial shear waves are Tkachenko modes [3, 4, 5, 6].

### v4

# Anomalous dispersion of shear waves in dipolar supersolids

@@ Abstract
Dipolar supersolids—quantum states that are simultaneously superfluid and solid—have had their superfluid nature rigorously tested, while its solid nature remains uncharted. Arguably, the defining characteristic of a solid is the existence of elastic shear waves. In this work, we investigate transverse wave packet propagation in dipolar supersolids with triangular and honeycomb structure. Remarkably, the honeycomb supersolid displays anomalous dispersion, supporting waves traveling faster than the transverse speed of sound. While the distinction between anomalous and normal dispersion appears subtle, it has a profound influence on both the transport and scattering properties. For both supersolid phases, we calculate the shear modulus, a key parameter that quantifies the material’s rigidity. Our findings are pertinent to current experimental efforts scrutinizing the fundamental properties of supersolids.
@@
Here, we propose a protocol to excite shear waves in two-dimensional dipolar supersolids, and measure their resulting propagation velocities. We investigate this protocol for triangular and honeycomb lattice structures, which are ground state solutions for a system of ultracold dipolar atoms under planar confinement [41, 42, 43]. Our results reveal distinct dispersive phenomena between the triangular and honeycomb phases, which we probe through the current density dynamic structure factor (DSF) and the Bogoliubov de-Gennes (BdG) method. These investigations show that the honeycomb structure can support a refractive index smaller than unity, thus supporting shear waves traveling faster than the transverse speed of sound. The idea of such a sound mode in classical wave theory can be traced back to Sommerfeld and Brillouin [44, 45, 46, 47], while in the quantum realm, it is analogous to the anomalous phonon dispersion relation observed in superfluid helium-4 [48, 49, 50, 23]. We speculate that this anomalous behavior is a signature of an incipient transverse branch instability in the honeycomb supersolid as recently reported by Blakie [51]. Finally, we find the relationship between the wave packet propagation velocities and the transverse speed of sound, and estimate the shear modulus of dipolar supersolids.
@@
where the dimensionally reduced wave function $\psi(\bm{\rho},t)$ is defined over the infinite plane, $\bm{\rho}=(x,y)$. The 2D interaction and quantum fluctuation terms, $\Phi_{\sigma}$ and $g_{\text{QF}}$, depend on the variational width $\sigma$ in the perpendicular ($z$) direction. We fix the areal number density $n_{\text{2D}}$, and perpendicular harmonic trap with frequency $\omega_{z}=2\pi\times 72.4\mathrm{\,Hz}$. This model and our numerical methods are detailed in the Supplemental Material [57].
@@
where $u_{\nu}$ and $v_{\nu}$ are the Bogoliubov modes with band index $\nu$ which obey the translational symmetry of the supersolid state. By substituting the perturbed wavefunction into Eq. (2) and linearizing, we obtain the familiar Bogoliubov-de Gennes (BdG) equations [67, 68] (see the Supplemental Material for details [57]). We solve the BdG equations using a Bloch wave expansion of the Bogoliubov modes to obtain the frequencies, $\omega_{\nu}(\bm{k})$, and overlay the three lowest bands for comparison in Fig. 1(c). Note that the near-identical shape of the superfluid and transverse modes is a coincidence of the chosen parameters, and their speeds of sound are distinct in other regions of parameter space [68]. We can extract three characteristic velocities from the BdG results for the transverse branch: 1) the transverse speed of sound $c_{t}=\lim_{k_{y}\rightarrow 0}\omega_{t}/k_{y}$, 2) the phase velocity $v_{p}(k_{y})=\omega_{t}/k_{y}$, which defines the velocity of a wave with given $k_{y}$, and 3) the group velocity, $v_{g}(k_{y})=\partial\omega_{t}/\partial k_{y}$, which defines the velocity of the wave packet envelope and thus the velocity of energy propagation in the medium  [^2] We extract these quantities from the BdG calculations by fitting the transverse branch with a fourth-order polynomial.. For the triangular supersolid, both $v_{g}$ and $v_{p}$ monotonically decrease from $c_{t}$ with increasing wave number, and subsequently the refractive index, $n=c_{t}/v_{p}$, increases from unity with $k_{y}$. This behavior is referred to as normal dispersion and is a hallmark of conventional crystalline solids [70].
@@
Anomalous dispersion of shear waves in honeycomb supersolids—Applying the same protocol to the honeycomb supersolid [Fig. 2(a)], we observe radically distinct behavior. For the triangular supersolid, the wave packet edge propagates at the speed of sound, where $\omega=0$, as indicated by an enduring rightward COM displacement (colored red in Fig. 1(b)). In contrast, the wave packet edge in the honeycomb supersolid oscillates between left and right displacement [Fig. 2(b)]. Thus, the fastest velocity in the excitation has $\omega>0$, and exceeds the speed of sound.
We can pinpoint the origin of this unexpected finding from the current density DSF, the BdG dispersion and the corresponding characteristic velocities, as shown in Figs. 2(c)-(d)  [^3] Note that as the perturbation dislocates the high density region, we excite many branches, but still dominantly transverse crystal modes.. The transverse branch displays a subtle upward concavity for intermediate wave vector ($0.2\lesssim k_{y}L/\pi\lesssim 0.4$), supporting $v_{g}$ and $v_{p}$ which exceed $c_{t}$, and a sub-unity refractive index. This is known as anomalous dispersion and has significant implications for transverse phonon decay. The superlinear dispersion allows for Beliaev damping of transverse phonons, where one phonon spontaneously decays into two [72, 73]. Though rare, similar phenomena has been observed in other systems such as crystalline materials [74, 75], Wigner and Coulomb crystals [76, 77, 78] and superfluid helium-4 [49], the latter being the subject of thirty years of intense experimental and theoretical debate [23].
The origin of the anomalous behavior can be understood through a simple physical argument, though this will be substantiated in a subsequent work. In general, the dispersion relation for a gapless branch must have the form $\omega(k)=Ak+Bk^{2}+\mathcal{O}(k^{3})$ for momentum $k$. A phonon instability occurs when the linear part becomes negative, crossing $A=0$. However, in a crystalline system at the Brillouin zone (BZ) edge $\text{d}\omega(k)/\text{d}k=0$. Therefore, near to the shear wave (transverse phonon) instability the band goes as $k^{2}$ for low momenta and flattens at the BZ edge. These criteria set “boundary conditions” for the band and lead to anomalous behavior for intermediate $k$. We note that this implies an enhancement of the anomalous behavior close to the shear wave instability [51], however, the honeycomb state is not the ground state in that region, and it is not be possible to dynamically probe this in an experiment without entering the stripe regime. This was shown in the top row of Fig. 3 in the very recent work of Ref. [51]. Our results show that this behavior persists in an experimentally relevant parameter regime where the honeycomb state is indeed the ground state. To our knowledge, the honeycomb supersolid constitutes the first quantum solid to display anomalous dispersion, the main result of this Letter.
@@
Consider the experimental implementation of this protocol for the triangular supersolid. The edge velocity matches closely with our extracted $c_{t}$ [Fig. 3(a1)], however due to fluctuations such as temperature effects, tracking the edge of the shear wave may be below the fidelity of experiments. Furthermore, in finite size systems the propagation velocity can be greatly reduced at the boundary of the system, as we show in the Supplemental Material [57]. Thus, experimental estimates may drastically underestimate the shear modulus. At much larger densities than shown here, the transverse branch becomes soft, and the shear wave velocity tends to zero. However, this is in the parameter regime where the stripe phase is the ground state solution [42].
The honeycomb state in Fig. 3(a2) shows a propagation velocity exceeding $c_{t}$ across the parameter space, however this effect is diminished for increasing densities as indicated by the decreasing width of the shaded region. This transition from anomalous to normal dispersion is reminiscent of the behavior seen in superfluid helium-4 as a function of pressure [23]. The anomalous dispersion is maximized at low densities, close to the honeycomb-to-stripe transition, and may be related to the hybridization of modes across this structural phase transition. Refs. [9, 79, 80, 81, 68] define Eq. (5) using the transverse speed of sound, replacing $v_{s}\rightarrow c_{t}$. The prediction of anomalously dispersive transverse waves calls this definition into question. Instead, the stiffness is a $k$-dependent quantity, and the shear modulus of the honeycomb supersolid can exceed the one calculated from the speed of sound.
Conclusions—We have demonstrated that dipolar supersolids with two-dimensional crystalline structure are true solids displaying a finite shear modulus. However, the triangular and honeycomb states exhibit shear wave phenomena of a diverging nature. The triangular lattice structure demonstrates normal dispersion, and this modulated superfluid state behaves as an archetypal solid. In contrast, supersolids with honeycomb lattice structure show anomalous dispersion, which supports shear wave propagation exceeding the transverse speed of sound and opens intriguing avenues for future theoretical and experimental exploration. We expect a rich phenomenology to arise in this phase from the novel scattering and spontaneous decay channels of anomalously dispersive transverse phonons.
@@
Our work opens up new avenues for non-invasive measurements of dipolar supersolids. In future work, we will explore how defects in the supersolid affect the propagation of shear waves, bringing the ideas of shear wave elastography to the quantum realm. A supersolid can host both solid and superfluid defects, as holes in the crystal structure and quantized vortices, respectively. These superfluid vortices lie in the low-density regions, and are not possible to image directly [38, 96]. How these defects distort and interact with shear waves [97] may provide an indirect measurement of their presence in future dipolar supersolid experiments. Our findings may also provide insights into the propagation of shear waves in the supersolid phase of a neutron star following a starquake or glitch [29].